Do=1500-10Po+4Y+Pa
So=800+2Po
Find demand equation for oranges in terms of price for oranges( Po),when Y is ₹500 and Pa is ₹60.
Find equilibrium price and quantity.
What is the excess demand or supply if government sets the price ₹250.
1
Expert's answer
2020-09-07T10:49:55-0400

The general demand equation is '

"Qd=a-bP"

Therefore the demand equation for oranges in terms of price is

"Do=1500-10Po+4Y+Pa\ where,\newline Y=500 rupees\newline Pa=60 rupees"

Substituting values in the equation

"Do=1500-10Po+2000+60\newline Do=3560-10Po"


Equilibrium price and quantity

Equilibrium price is a market price where the amount of goods demanded is equal to the amount of goods supplied in the market.

The quantity demand equation of oranges

,"Do=3560-10Po" where,

Do is the quantity of oranges a consumer wants to buy

Po is the price of oranges

The quantity supplied equation,

"So=800+2Po" where

So is the quantity of oranges the producer will supply

Equilibrium price is at the point where demands equals the supply.

"Do=So"

Substituting the values

"3560-10Po=800+2Po,,,solving \ with\ algebra\newline 2760=12Po\newline Po=230"

"equlibrium\ price = 230"

To get equilibrium quantity substitute the price value in the demand equation i.e

"Do=3560-10(230)\newline Do=1260"

"equlibrium \ quanitity =1260"


If the government sets the price to be 250 rupees the supply will be excess.

"Do=3560-10(250)=1060\newline So=800+2(250)=1300\newline Quanitity\ supplied\ will \ be\ excess"